# Using matrix-product states for time-series machine learning

### v1

@@ II.4 Classification
Since all classes are assumed to have the same prior probability $p(l)$, the probability of each class, given an unlabeled time series, $p(l\mid\mathbf{x})$, is directly proportional to its non-normalized likelihood $p(\mathbf{x}\mid l)$ i.e., $p(l\mid\mathbf{x})\propto p(\mathbf{x}\mid l)$ according to Bayes’ rule. Therefore, the predicted class can be determined directly from the model output by selecting the index $l$ corresponding to the largest entry in $f^{l}(\mathbf{x})$:
@@
In this section, we have introduced an MPS-based algorithm for learning a generative distribution underlying an observed time-series dataset. By leveraging the properties of the MPS to encode a tractable representation of complex joint distributions, it is possible to sample from and do inference on those distributions, unlocking the potential to tackle a wide range of important time-series analysis problems. Crucially, our approach allows us to perform these tasks, including imputation and classification, under a unified and statistically grounded framework.
@@ III Numerical Experiments
In Sec. II above, we introduced a practical algorithm for learning an underlying probabilistic model from a time-series dataset, detailing how the encoded generative distribution can then be used as the basis of new MPS-based algorithms for time-series analysis tasks. While existing methods have been developed to address time-series ML tasks, such as classification and imputation independently, our approach is the first to tackle these problems within a common modeling framework by leveraging a tractable form of the complex joint distribution underlying time series. In this section, we aim to investigate the utility of MPSTime in a range of important data-driven problems involving synthetic and real-world time-series datasets. First, in Sec. III.1, we detail the construction of the datasets we investigate in our numerical experiments. Then, in Sec. III.2, we empirically validate our framework for MPS-based imputation on a synthetic dataset, showing that MPSTime can successfully learn the underlying joint distribution directly from time-series data.
@@ III.1 Time-series datasets
To evaluate our MPS-based algorithm for time-series learning and inference, we investigated several synthetic and real-world datasets through a series of carefully constructed tasks. Here, we detail the pre-processing steps necessary to derive our classification and imputation tasks from publicly available time-series datasets. Table 1 provides a high-level overview of the datasets used in our experiments.
@@ III.1.1 Synthetic datasets
We first aimed to validate the ability of MPSTime to infer distributions over time-series data by simulating time series from an analytic generative model. To this end, we investigated two synthetic test cases in which we simulated time series of length $T=100$ samples from a ‘noisy trendy sinusoid’ (NTS) model:
@@
where $x_{t}$ is the observed value at time $t\in[0,T]$, $\tau$ is the period, $m$ is the slope of a linear trend, $\psi\in[0,2\pi)$ is the phase offset, $\sigma$ is the noise scale, and $n_{t}\sim\mathcal{N}(0,1)$ are i.i.d. normally distributed random variables. A time-series dataset can be constructed by repeatedly generating time series from Eq. (22). The diversity of the time-series dataset (and thus complexity of the joint distribution over time series) can be controlled via the parameters of Eq. 22 that are fixed for all time series in the dataset, versus the parameters that are allowed to vary across the dataset. Here we analyzed six NTS datasets of differing complexity, generated through different such choices, as detailed below:
@@
    Simple settings: In our first test case, we fixed the period ($\tau=20$) and trend ($m=3$), while allowing the phase offset $\psi$ to vary across the time-series dataset. This was achieved by sampling a value for $\psi$ for each time series as a random sample from a uniform distribution, as $\psi\sim U(0,2\pi)$. We generated two training datasets containing $300$ time series each: one with: (i) NTS1, with low noise ($\sigma=0.1$); and (ii) NTS2, with moderate noise ($\sigma=0.2$). For imputation, we then generated a test dataset containing 200 time series on which to evaluate model performance.
@@
    Challenging settings: To simulate time series exhibiting richer dynamical variation, and therefore, more complex data distributions, we systematically expanded the parameter space of the NTS model. Fixing the noise scale at $\sigma=0.1$, we constructed four datasets with progressively more complex joint distributions by varying the trend $m$ and period $\tau$ parameters: (i) NTS3, with a single trend ($m=20$) and three periods ($\tau\in\{20,30,40\}$); (ii) NTS4, with three trends ($m\in\{-3,0,3\}$) and a single period ($\tau=20$); (iii) NTS5 with two trends ($m\in\{-3,3\}$) and two periods ($\tau\in\{20,40\}$); and (iv) NTS6, with three trends ($m\in\{-3,0,3\}$) and three periods ($\tau\in\{20,30,40\}$). In all cases, the phase offset $\psi$ was drawn randomly from a uniform distribution $\psi\sim U(0,2\pi)$, and the free parameter(s) were selected from randomly from the set of allowed values. In each case, we generated a training dataset of length $600$ time series to adequately sample the larger model parameter space. In each case, model performance was evaluated on a subsequent test set of $300$ generated time series.
@@ III.1.2 Real-world datasets
We selected three real-world time-series datasets to analyze, representative of three key application domains: (i) medicine, (ii) energy, and (iii) astronomy:
@@
    ECG: Our medical dataset, ECG200 [51], is a two-class dataset of time series representing electrocardiogram (ECG) traces of single heart beats, labeled according to whether the signal originated from a healthy or diseased (myocardial infarction) patient. The problem is well studied in the time-series classification (TSC) literature. The dataset comprises 100 training instances and 100 test instances, where each time series contains $T=96$ samples. Of the $200$ total time series, $133$ ($66.5\%$) are labeled as ‘healthy’, and the remaining 67 ($33.5\%$) are labeled as ‘abnormal’. For classification and imputation tasks, we used the same open dataset published in the UCR time-series classification archive [52]. However, for imputation, we discarded label information and effectively treat all instances as belonging to the same data distribution (i.e., a single ‘class’).
@@
    Power Demand: ItalyPowerDemand [53] is a two-class dataset of time series corresponding to one day of electrical power demand in a small Italian city, sampled at 1-hour intervals, with each time series assigned to either ‘winter’ (recorded between October and March) or ‘summer’ (recorded between April and September). The dataset is split into $67$ training instances and $1029$ test instances, each of length $T=24$ samples. Of the total $1096$ time series, 547 (49.9%) are labeled as ‘winter’ (class 0) and the remaining 549 (50.1%) are labeled as ‘summer’ (class 1). For classification and imputation, we use the same open dataset published in the UCR time-series classification archive [52], omitting class labels for the latter task.
@@
    Astronomy: The KeplerLightCurves [54] dataset comprises $1319$ light curves from NASA’s Kepler mission. Individual time series correspond to stellar brightness measurements from a single star, sampled every $30$ minutes over a three-month quarter (‘Quarter 9’), yielding length $T=4767$ sample instances. For our classification task, we selected two of the seven classes that were particularly challenging to distinguish by eye: (i) non-variable stars, and (ii) $\delta$ Scuti stars. We then applied random under-sampling to the majority class to yield a balanced class distribution (201 non-variable stars and 201 $\delta$ Scuti stars). Instances were randomly assigned to either the train or test set using an 80/20 train/test ratio, and the first 100 samples of each time series were retained for classification.
    For the imputation task, we selected two classes from the KeplerLightCurves dataset: (i) RR Lyrae variable (25 time series), and (ii) non variable (25 time series). Unlike Power Demand and ECG, here we focused on the more typical imputation setting involving a single time-series instance. To construct the imputation datasets, each $T=4767$ sample time series was truncated to $4700$ samples and split into 47 non-overlapping windows of $T=100$ samples each. Windows containing true missing data gaps were discarded, and 80% of the remaining ‘clean’ windows (43 windows) were randomly allocated to the training set, with 20% (2 windows) allocated to the test set.
Table 1: Summary of all simulated and real-world time-series datasets analyzed here. Datasets are separated into those used for classification (labeled ‘C’) and imputation (labeled ‘I’) tasks, see Sec. III.1 for details.
Dataset Train Size Test Size Length Classes
ECG (C) 100 (50%) 100 (50%) 96 2
Power Demand (C) 67 (6%) 1029 (94%) 24 2
Astronomy (C) 321 (80%) 81 (20%) 100 2
NTS1–2 (I)§ 300 (60%) 200 (40%) 100 1
NTS3–6 (I) 600 (67%) 300 (33%) 100 1
ECG (I) 100 (50%) 100 (50%) 96 1
Power Demand (I) 67 (6%) 1029 (94%) 24 1
Astronomy (I) 43 (96%) 2 (4%) 100 1
§NTS1–2 refers to datasets NTS1 and NTS2. For Astronomy, we use ‘dataset’ to refer to the collection of fixed-length windows extracted from a single time-series instance. A total of 50 such datasets (i.e., 50 windowed time series) were evaluated.
@@ III.2 Time-series imputation
So far, we have presented a practical and statistically grounded framework for MPS-based probabilistic modeling of time-series data. While the theoretical foundations are well-established, the practical efficacy of our approach – particularly the capacity to learn complex joint distributions for time-series ML problems – remains to be demonstrated. To bridge this gap between theory and practice, we now turn to an experimental validation of our approach, demonstrating first the ability of MPS to model complex joint distributions, and second, leveraging these learned distributions to perform an accurate imputation of missing values.
We begin by validating our approach on a relatively simple joint distribution of trendy sinusoids, providing a compelling proof-of-concept for the effectiveness of MPS-based imputation in a controlled setting. By systematically varying the parameters of our synthetic data generative model, we then show that a sufficiently expressive MPS can capture increasingly complex joint distributions of time-series data. Moving beyond synthetic examples, we apply our approach to impute missing values in three real-world datasets with unknown generative processes.
In order to demonstrate the performance of our MPS-based time-series imputation approach under varying degrees of data loss, we applied our method to several synthetic and real-world datasets. To simulate realistic scenarios often encountered in practical sensor deployments – such as sensor drop-outs due to misplacement, hardware failure, or data-downlinks – we focused on the contiguous missing data setting, where blocks of consecutive time-series observations were removed from test set instances. We note that if missing data are unmeasured future values, the imputation task can be considered a type of forecasting problem. Imputation performance was evaluated using the mean absolute error (MAE) defined as:
@@
where $y_{i}$ is the actual missing value for the $i$-th data point, $\hat{y}_{i}$ is the imputed value for the $i$-th data point, and $n$ is the total number of missing values. A smaller MAE implies a higher accuracy.
@@
To evaluate the imputation capabilities of our method, we varied the proportion of missing data points in each test instance by increasing the percentage of missing data from 5% to 95% of all samples. We then compared the imputation performance of our approach to several baselines, each selected to encompass a diverse range of approaches in the time-series imputation literature: (i) 1-Nearest Neighbor Imputation (1-NNI); (ii) Centroid Decomposition Recovery (CDRec); (iii) Bidirectional Recurrent Imputation for Time Series (BRITS); and, (iv) Conditional Score-based Diffusion Imputation (CSDI).
@@
The classical baseline, 1-NNI [46], imputes missing data points by identifying the nearest neighbor in the training set (based on a Euclidean distance of observed samples), then substitutes the missing values with those from the neighboring time series. CDRec [55] decomposes a time-series matrix (i.e., an $N\times T$ matrix of $N$ time-series instances, each with fixed length $T$) into centroid patterns that capture recurring temporal behaviors, enabling the recovery of missing values through pattern-based reconstruction. The generative modeling baseline, CSDI, [56] uses a neural network-based diffusion model to gradually convert random noise into plausible imputed data points that are consistent with the data distribution conditioned on observed values. Finally, BRITS [57] is a popular deep learning method which leverages recurrent neural networks (RNNs) to model temporal dependencies in both forward and backward directions. While CDRec, CSDI, and BRITS were initially intended for multivariate time-series imputation, CDRec is one of the best performing algorithms in ‘total blackout’ scenarios [58], while CSDI is the best performing generative model-based imputation algorithm [42, 59], and most closely matches the distribution-learning approach of MPSTime.
@@
Figure 4: Time-series imputation of synthesized dataset. We compare MPSTime’s (solid line) imputation performance against the 1-nearest neighbor imputation (1-NNI) baseline (dotted line) on synthetic datasets of phase-randomized, noisy trendy sinusoids generated by the model defined in Eq. (22). The test-set mean absolute error (MAE) between the imputed values and ground-truth (unobserved) values is reported across varying percentages of missing data. (a) Simple setting: MPS with physical dimension $d=12$ and three different bond dimensions $\chi_{\rm{max}}=20,30,40$ (blue, orange, and green, respectively) trained on datasets with fixed trend $m$, fixed period $\tau$, and two noise levels $\sigma=0.1$ (NTS1, circle marker) and $\sigma=0.2$ (NTS2, diamond marker). (b) Challenging setting: MPS with $d=20$ and $\chi_{\rm{max}}=150$ trained on four datasets (NTS3-6) with increasingly complex underlying joint distributions, constructed by varying the trends $m$ and/or periods $\tau$ of noise-corrupted, phase-randomized sinusoids with $\sigma=0.1$, for the model defined in Eq. (22). In both (a) and (b), we show representative time-series examples of the MPS-imputed values on unseen (‘test’) time series generated from the same model [NTS2 dataset with $\chi_{\rm max}=30$ for (a) and NTS6 dataset with $\chi_{\rm max}=150$ for (b)] in the panels (ii)–(vii). The red shading represents the uncertainty, which here is quantified using the weighted median absolute deviation (WMAD).

### v2

@@ II.4 Time-series classification algorithm
As all classes are assumed to have the same prior probability $p(l)$, then the probability of each class given an unlabeled time series, $p(l\mid\mathbf{x})$, is directly proportional to its non-normalized likelihood $p(\mathbf{x}\mid l)$ i.e., $p(l\mid\mathbf{x})\propto p(\mathbf{x}\mid l)$ (cf. Bayes’ rule [54]). Therefore, the predicted class can be determined directly from the model output by selecting the index $l$ corresponding to the largest entry in $f^{l}(\mathbf{x})$:
@@
In this section, we have introduced MPSTime, an MPS-based algorithm for estimating the joint distribution $p(\mathbf{x})$ underlying a dataset of time series. By leveraging the properties of the MPS to encode an approximation of complex joint distributions, it is possible to sample from and do inference on those distributions, allowing it to form the basis for novel time-series machine-learning problems including imputation and classification. Crucially, while existing algorithms often address different problem classes like classification and imputation using specifically tailored algorithms, MPSTime can tackle them within a unified time-series modeling framework.
@@ III Numerical Experiments
To validate the proposed MPSTime framework developed above, in this section we conduct a range of numerical experiments aimed to demonstrate its ability to approximate time-series joint distributions, and apply this structure to practical time-series ML tasks. We begin in Sec. III.1 by detailing the construction of the datasets comprising the tasks. Then, in Sec. III.2, we empirically validate our framework for MPS-based imputation on these datasets, showing that MPSTime can successfully learn the underlying joint distribution $p(\mathbf{x})$ directly from time-series data.
@@ III.1 Time-series datasets
To validate our MPS-based algorithm for time-series machine learning (ML), we investigated several synthetic and real-world datasets. In this work, we focus on the tasks of time-series imputation and classification, selected to encompass the breadth of application of MPS-based inference for time series. Here, we detail the process to derive the datasets from publicly available repositories. Table 1 provides a summary of the datasets used in our experiments.
@@ III.1.1 Synthetic datasets
We first aimed to validate the ability of MPSTime to infer distributions over time-series data by simulating time-series instances from an analytic generative model. To achieve this, we generated time series of length $T=100$ samples from a ‘noisy trendy sinusoid’ (NTS) model:
@@
where $x_{t}$ is the observed value at time $t\in[0,T]$, $\tau$ is the period of the sinusoid, $m$ is the slope of a constant linear trend, $\psi\in[0,2\pi)$ is the phase offset, $\sigma$ is the noise scale, and $n_{t}\sim\mathcal{N}(0,1)$ are i.i.d. normally distributed random variables. Time-series datasets were then constructed by repeatedly generating time-series instances as realizations from Eq. (22). The diversity of the time-series dataset (and thus complexity of the joint distribution over time series) can be controlled via the parameters of Eq. (22) that are fixed for all time series in the dataset, versus the parameters that are allowed to vary across the dataset. We constructed five NTS datasets of differing complexity, generated through different numbers and degrees of parametric freedoms, as detailed below:
@@
    Simple setting: For the simplest dataset, NTS1, we fixed the period ($\tau=20$), trend ($m=3$) and noise scale ($\sigma=0.1$) in Eq. (22), while allowing the phase $\psi$ to vary across the time-series dataset. This variation was achieved by sampling a value for $\psi$ for each time series as a random sample from a uniform distribution, as $\psi\sim U(0,2\pi)$. We generated a training set containing $300$ time series, and a test set of $200$ time series from the same generative model.
@@
    Challenging settings: To simulate time series exhibiting richer dynamical variation, and therefore more complex data distributions, we systematically expanded the parameter space of the NTS model. Fixing the noise scale at $\sigma=0.1$, we constructed four datasets with progressively more complex data distributions by varying the trend $m$ and period $\tau$ parameters: (i) NTS2, with a single trend ($m=3.0$) and three periods ($\tau\in\{20,30,40\}$); (ii) NTS3, with three trends ($m\in\{-3,0,3\}$) and a single period ($\tau=20$); (iii) NTS4 with two trends ($m\in\{-3,3\}$) and two periods ($\tau\in\{20,40\}$); and (iv) NTS5, with three trends ($m\in\{-3,0,3\}$) and three periods ($\tau\in\{20,30,40\}$). For all datasets, the phase offset $\psi$ was drawn randomly from a uniform distribution $\psi\sim U(0,2\pi)$, and the free parameter(s) were selected randomly from the set of allowed values. For NTS2, NTS3, and NTS4, we generated a training dataset containing $400$ time series (and $500$ for the most complex NTS5 model) to densely sample the larger model parameter space, and all models were evaluated on a test set of $200$ time series (generated from the same model and parameter settings).
@@ III.1.2 Real-world datasets
To evaluate the performance of MPSTime on data for which the underlying generative process is unknown, we selected three real-world time-series datasets to analyze, representative of three key application domains: (i) medicine, (ii) energy, and (iii) astronomy:
@@
    ECG: This medical dataset, labeled as ECG200 in the UCR TSC archive [55], is a two-class dataset of time series representing electrocardiogram (ECG) traces of single heart beats, labeled according to whether the signal originated from a healthy or diseased (myocardial infarction) patient [56]. The dataset comprises 100 training instances and 100 test instances, where each time series contains $T=96$ samples. Of the $200$ total time series, $133$ ($66.5\%$) are labeled as ‘healthy’, and the remaining 67 ($33.5\%$) are labeled as ‘abnormal’. For classification and imputation tasks, we used the same open dataset published in the UCR time-series classification archive [55].
@@
    Power Demand: The Power Demand dataset, labeled as ItalyPowerDemand in the UCR repository [55], is a two-class dataset of time series corresponding to one day of electrical power demand in a small Italian city, sampled at 1 h intervals [57]. Each time series is assigned to either ‘winter’ (recorded between the months of October and March) or ‘summer’ (recorded between the months of April and September). The dataset is split into $67$ training instances and $1029$ test instances, each of length $T=24$ samples. Of the total $1096$ time series, 547 (49.9%) are labeled as ‘winter’ and the remaining 549 (50.1%) are labeled as ‘summer’. For classification and imputation, we use the same open dataset published in the UCR time-series classification archive [55].
@@
    Astronomy: The Astronomy dataset is derived from the UCR TSC archive KeplerLightCurves dataset [55] and comprises $1319$ light curves from NASA’s Kepler mission [58]. Individual time series correspond to stellar brightness measurements from a single star, sampled every $30$ minutes over a three-month quarter (‘Quarter 9’), yielding length $T=4767$ sample instances. For our classification task, we selected two of the seven classes that were particularly challenging to distinguish by eye which were (i) non-variables, and (ii) $\delta$ Scuti stars. We then applied random under-sampling to the majority class ($\delta$ Scuti stars, originally containing 411 instances) to yield a balanced class distribution (201 non-variable stars and 201 $\delta$ Scuti stars). Instances were randomly assigned to either the train or test set using an 80/20 train/test ratio, and each time-series instance was truncated to its first 100 samples for classification.
    For the imputation task, we selected two different classes from the KeplerLightCurves dataset: (i) RR Lyrae stars; and (ii) $\gamma$ Doradus stars. Unlike the Power Demand and ECG datasets, here we focused on the typical setting in which imputation is performed within a single time-series instance by using observed values to infer missing values. To construct the imputation datasets, we randomly selected five time-series instances from each class. These instances, each of length $T=4767$ samples was truncated to $4700$ samples and segmented into 47 non-overlapping windows of $T=100$ samples. Windows containing true missing data gaps were discarded, and 80% of the remaining ‘clean’ windows (43 windows) were randomly allocated to the training set, with 20% (2 windows) allocated to the test set.
Table 1: Summary of all simulated and real-world time-series datasets analyzed. Datasets are categorized into those used for classification (labeled ‘C’) and imputation (labeled ‘I’) tasks, see Sec. III.1 for details.
Dataset Train Size Test Size Length
ECG (C) 100 (50%) 100 (50%) 96
Power Demand (C) 67 (6%) 1029 (94%) 24
Astronomy (C) 321 (80%) 81 (20%) 100
NTS1 (I) 300 (60%) 200 (40%) 100
NTS2–4 (I)§ 400 (67%) 200 (33%) 100
NTS5 (I) 500 (71%) 200 (29%) 100
ECG (I) 100 (50%) 100 (50%) 96
Power Demand (I) 67 (6%) 1029 (94%) 24
Astronomy (I) 43 (96%) 2 (4%) 100
§NTS2–4 refers to datasets NTS2, NTS3, and NTS4.
For consistency, we use ‘dataset’ to refer to the collection of fixed-length windows extracted from a single time-series instance in the Astronomy (I) dataset. A total of 10 such datasets (i.e., 10 independent and windowed time-series instances) were evaluated.
@@ III.2 Time-series imputation
The recovery of missing time-series values via imputation represents a key challenge in time-series analysis, as real-world data collection rarely occurs under ideal conditions. Sensor dropouts due to malfunction, operational constraints, or other disruptions, can give rise to contiguous blocks of unobserved time-series values. Robust imputation methods that accurately reconstruct missing values, while preserving the underlying temporal structure of the data, are essential for reliable analysis and inference on such datasets.
In this work, we specifically focus on univariate time-series imputation in a dataset-level setting, where a model is trained on a collection of $N$ ‘clean’ (non-corrupted) time-series instances of fixed length $T$. At test time, imputation is performed on unseen, partially corrupted instances of the same length $T$ by conditioning on the observed values, and reconstructing unobserved values from the joint distribution learned on the training set. We note that this setup differs from the typical univariate imputation setting, where reconstruction is performed within a single, arbitrarily long time series, and using only its own observed values [59, 60].
We now turn to an experimental validation of our MPS-based approach to imputation, demonstrating first the ability of MPSTime to model complex joint distributions, and second, leveraging these learned distributions to perform an accurate imputation of missing values. We begin by validating our approach on a relatively simple joint distribution of trendy sinusoids. By systematically varying the parameters of our synthetic data generative model, we then show that a sufficiently expressive MPS can capture increasingly complex joint distributions of time-series data. Moving beyond synthetic examples, we apply our approach to impute missing values in three real-world datasets with unknown generative processes.
In order to demonstrate the robustness of our MPS-based approach under varying degrees of data loss, we varied the proportion of missing data points in each imputation task by increasing the percentage of missing data from 5% to 95% of all samples. In this work, we specifically focused on the contiguous missing data setting, where blocks of consecutive time-series observations were removed from test set instances. Note that if missing data are unmeasured future values, such an imputation task could be considered a type of forecasting problem. Imputation performance was evaluated in the original data domain using the mean absolute error (MAE) defined as:
@@
where $y_{i}$ is the actual missing value for the $i$-th data point, $\hat{y}_{i}$ is the imputed value for the $i$-th data point, and $n$ is the number of missing values.
@@
We compared the imputation performance of our approach to several benchmarks, each selected to encompass a diverse range of approaches in the time-series imputation literature including: (i) 1-Nearest Neighbor Imputation (1-NNI); (ii) Centroid Decomposition Recovery (CDRec); (iii) Bidirectional Recurrent Imputation for Time Series (BRITS-I); and (iv) Conditional Score-based Diffusion Imputation (CSDI).
@@
The simple benchmark, 1-NNI [49], imputes missing data points by identifying the nearest neighbor in the training set (based on a Euclidean distance of observed samples), then substitutes the missing values with those from the neighboring time series. CDRec [61] decomposes a time-series matrix (i.e., an $N\times T$ matrix of $N$ time-series instances, each with fixed length $T$) into centroid patterns that capture recurring temporal behaviors, enabling the recovery of missing values through pattern-based reconstruction. The generative modeling benchmark, CSDI [62] uses a neural network-based diffusion model to gradually convert random noise into plausible imputed data points that are consistent with the data distribution conditioned on observed values. Finally, BRITS-I [63] is a popular deep-learning algorithm which uses recurrent neural networks (RNNs) to model temporal dependencies in both forward and backward directions. While CDRec and CSDI were initially intended for multivariate time-series imputation, CDRec is one of the best performing algorithms in ‘total blackout’ scenarios [64], and CSDI is the best performing generative model-based imputation algorithm [45, 65], which most closely matches the distribution-learning approach of MPSTime.
@@
Figure 4: Matrix-Product State (MPS)-based time-series imputation of synthesized datasets. We compare MPSTime (solid line) against the 1-nearest-neighbor imputation (1-NNI) benchmark (dashed black line) on synthetic datasets of phase-randomized, noisy trendy sinusoids (NTS) generated by the model defined in Eq. (22). The test set mean absolute error (MAE) between the imputed values and ground-truth (unobserved) values is reported across varying percentages of missing data. (a) Simple setting: MPSTime with three different bond dimensions $\chi_{\rm{max}}=20,30,40$ (blue, orange, and green, respectively) is trained on a dataset exhibiting limited dynamical variation and a correspondingly simpler joint distribution with fixed trend $m$, fixed period $\tau$, and noise level $\sigma=0.1$. (b) Challenging setting: MPSTime with parameters determined by hyperparameter tuning trained on four datasets NTS2–5 (blue, orange, green, red, respectively) that exhibit richer dynamical variability and increasingly complex underlying joint distributions, constructed by varying the trends $m$ and/or periods $\tau$ of noise-corrupted, phase-randomized sinusoids with $\sigma=0.1$, for the model defined in Eq. (22). The upper-right inset zooms in on missing data percentages from $5\%$ to $85\%$ to highlight performance differences across the datasets. In both (a) and (b), we show representative time-series examples of the MPSTime-imputed values on unseen (‘test’) time series generated from the same model [NTS1 dataset with $\chi_{\rm max}=40$ for (a) and NTS5 dataset with $\chi_{\rm max}=160$ for (b)] in the panels (ii)–(v). The red shading in representative time-series examples indicates the imputation uncertainty, which here is quantified using the weighted median absolute deviation (WMAD).